### v2

# Multi-Agent Taskforce Collaboration: Self-Correction of Compounding Errors in Long-Form Literature Review Generation

Abstract
Compounding error is critical in long-form literature review generation, where minor inaccuracies cascade and amplify across subsequent steps, severely compromising the faithfulness of the final output. To address this challenge, we propose the Multi-Agent Taskforce Collaboration (MATC) framework, which proactively mitigates errors by orchestrating LLM-based agents into three specialized taskforces: (1) an exploration taskforce that interleaves retrieval and outlining using a tree-based strategy to establish a grounded structure; (2) an exploitation taskforce that iteratively cycles between fact location and draft refinement to ensure evidential support; and (3) a feedback taskforce that leverages historical experience for self-correction before errors propagate. Experimental results show that MATC achieves state-of-the-art performance on existing benchmarks (AutoSurvey and SurveyEval), significantly outperforming strong baselines in both citation quality (e.g., +15.7% recall) and content quality. We further contribute TopSurvey, a new large-scale benchmark of 195 peer-reviewed survey topics, on which MATC maintains robust performance, demonstrating its generalizability.
Multi-Agent Taskforce Collaboration: Self-Correction of Compounding Errors in Long-Form Literature Review Generation
Zhi Zhang1, Yan Liu1thanks: Corresponding author., Zhejing Hu1, Gong Chen2, Sheng-hua Zhong3, Jiannong Cao1 1Department of Computing, The Hong Kong Polytechnic University, Hong Kong, China 2FireTorch Partners, Shenzhen, China 3College of Computer Science and Software Engineering, Shenzhen University, Shenzhen, China {zhi271.zhang, zhejing.hu}@connect.polyu.hk, {yan.liu, jiannong.cao}@polyu.edu.hk, heinz@clozzz.com, csshzhong@szu.edu.cn
1 Introduction
Literature reviews play an important role in scientific research by synthesizing existing knowledge to identify research gaps, establish theoretical frameworks, and guide future studies Ermel et al. (2021); Liu et al. (2023). Over recent decades, the rapid growth in the number of scientific publications has made it increasingly challenging for researchers to efficiently retrieve, organize, and synthesize relevant information in their fields Wang et al. (2024b). In response, automatic literature review systems that leverage advances in artificial intelligence have become increasingly valuable, as they help reduce manual workload, provide timely coverage of new research, and facilitate the identification of key trends and research gaps within the extensive literature Zhang et al. (2025); Wang et al. (2024b); Liu et al. (2023).
Early approaches primarily leverage multi-document summarization techniques to generate summaries from a set of reference papers Hoang and Kan (2010); Hu and Wan (2014); Erera et al. (2019), enabling the automatic creation of related work sections given a set of references as input. In recent years, large language models (LLMs) demonstrate impressive planning and reasoning abilities Wang et al. (2024a); Zhao et al. (2024); Zhong and Wang (2024); Nandy and Bandyopadhyay (2025); Zhang et al. (2025). This leads to a growing research area that employs LLMs as controllers to construct autonomous agents capable of effectively performing complex tasks Wang et al. (2024a). Building on these advances, recent state-of-the-art research, e.g., AutoSurvey Wang et al. (2024b) and SurveyX Liang et al. (2025), achieves remarkable success by designing LLM-based autonomous agents to automate the entire pipeline of literature review and generate long-form manuscripts.
Although these advances enable the automatic generation of human-like literature reviews with reduced manual effort, previous studies on multi-step models show that the inherent long-horizon workflow of such systems inevitably introduces the risk of compounding errors. In the workflow, small mistakes made at earlier steps can accumulate and propagate through subsequent steps, as errors that occur at a certain step will affect the next state Wang et al. (2025a). Without effective self-correction mechanisms, these deviations will further expand the error in later steps Han and Zhang (2023); Venkatraman et al. (2015); Hejna et al. (2023).
Figure 1: Illustration of the error accumulation phenomenon.
This paper investigates the phenomenon of error accumulation in automatic literature review generation. As a demonstration, we employ the state-of-the-art large language model GPT-4.1 OpenAI (2025) to conduct a literature review generation workflow on the topic of EEG (electroencephalogram)-based emotion recognition, with a target length of 4,000 words. As shown in Fig. 1, when the retrieval stage fails to identify relevant literature, the resulting errors can mislead the outlining process. An irrelevant outline may further cause the drafted manuscript to deviate from the intended topic. In long-form text generation, the increased complexity and the introduction of more dependent steps further exacerbate the compounding errors.
To tackle this challenge, we propose the Multi-Agent Taskforce Collaboration (MATC) framework. MATC consists of two types of autonomous agents: the manager agent and the executor agents. The manager agent orchestrates the workflow, while the executor agents each specialize in a distinct step of the process, such as literature searching, outline generation, fact localization, or manuscript drafting.
Inspired by organizational quality management Hackman and Wageman (1995), we propose to orchestrate agents organized as taskforces. A taskforce refers to a temporary group of specialized agents focused on specific subtasks Altier (1987); Aikin (1949); de Waard and Kramer (2008). We formulate three types of taskforces, each designed with specific error mitigation mechanisms.
Given a user’s instruction, the manager agent oversees the entire workflow and initiates the formation of different taskforces. To mitigate error accumulation between searching and outlining agents, the manager organizes an exploration taskforce employing a tree-based strategy. The taskforce begins with a broad overview and incrementally determines the literature and outline at each level, preventing errors in ungrounded outlines or biased retrieval. To mitigate error accumulation between locating and drafting agents, the manager organizes an exploitation taskforce employing an iterative cycle: the draft guides fact location from the literature, while the located facts inform draft refinement, preventing errors in unsupported claims or mismatched evidence. To mitigate errors within each agent’s workflow, the manager organizes feedback taskforces that maintain historical experience and implement dynamic checklists to guide self-correction before errors propagate to subsequent stages. Finally, the manager integrates the outputs of all taskforces to produce the final manuscript.
2 Related Work
Automating the summarization of research papers has long attracted interest. Early work mainly used multi-document summarization techniques. In 2010, Hoang et al. introduced automatic related work summarization, extracting sentences from multiple articles to form an extractive summary tailored to the target paper Hoang and Kan (2010). Lu et al. constructed Multi-XScience, a large-scale dataset for generating related work sections based on a paper’s abstract and cited articles, serving as a benchmark for later studies Lu et al. (2020). Chen et al. proposed a relation-aware multi-document encoder and a relation graph to capture cross-document dependencies for abstractive related work generation Chen et al. (2021). Xiao et al. developed PRIMERA, a pre-trained encoder-decoder model with novel objectives for linking and aggregating information across documents Xiao et al. (2022). Liu et al. proposed aligning sentences by category and using sparse transformers to organize information from many references Liu et al. (2023).
Figure 2: The overall framework of Multi-agent Taskforce Collaboration (MATC) for literature review generation.
Recently, advances in LLMs have greatly improved the scope and quality of automated literature reviews, enabling end-to-end systems to generate literature review papers. Wang et al. introduced AutoSurvey, a framework with retrieval, outline generation, parallel subsection drafting, integration, and rigorous evaluation Wang et al. (2024b). Shao et al. developed STORM, which models the pre-writing stage by discovering diverse perspectives, simulating expert conversations, and curating information to create an outline used to generate a full Wikipedia article Shao et al. (2024). Liang et al. designed SurveyX, an end-to-end solution for automated survey generation, covering online literature search, organization, and survey writing Liang et al. (2025). While these methods represent impressive progress, little research has addressed the risk of errors in generated outputs. Self-correction algorithms in LLMs are not suitable for the workflow of literature review generation Madaan et al. (2023); Shinn et al. (2023); Ferraz et al. (2024). As systems become more complex, error accumulation becomes a significant challenge.
3 Methodology
As illustrated in Fig. 2, given a user instruction $U$, the proposed Multi-Agent Taskforce Collaboration (MATC) framework orchestrates three taskforces to produce the final literature review output $Z$. In the following sections, we introduce each taskforce in detail.
3.1 Exploration Taskforce
The exploration taskforce is designed to determine the outline of the literature review and to retrieve relevant references for each section of the outline. Without prior knowledge of the relevant literature, the initial draft outline tends to be suboptimal. Conversely, without a determined direction, the literature retrieval process lacks focus. To mitigate compounding errors, we propose a tree-based strategy.
In detail, given a user instruction $U$, the manager $A_{M}$ initiates the exploration taskforce. The manager constructs a tree with the user instruction as the root node, representing the overall direction for exploring the field. This tree is expanded by identifying sub-directions of interest. For the root node, $A_{M}$ first invokes the searching agent to retrieve literature based on the user instruction, resulting in a set of relevant papers. The searching agent $A_{S}$ operates as follows:
$\{L^{(0)}_{1},L^{(0)}_{2},\dots,L^{(0)}_{I}\}=A_{S}(U)$ (1)
where $L_{i}^{(0)}$ denotes the $i$-th piece of literature retrieved for the root node at tree depth $d=0$. The number of literature is denoted by $I$.
Next, the manager assigns the outlining agent to determine sub-directions based on the titles and abstracts of the retrieved literature. The outlining agent $A_{O}$ operates as follows:
$\{O_{1}^{(0)},O_{2}^{(0)},\ldots,O_{J}^{(0)}\}=A_{O}(\{L^{(0)}_{i}\})$ (2)
where $O_{j}^{(0)}$ denotes the $j$-th sub-direction at the root level, and $J$ is the total number of sub-directions at tree depth $d=0$.
For each outline node $O_{j}^{(d)}$ at depth $d$, the manager decides whether further decomposition is needed according to the following criterion:
$\phi(O_{j}^{(d)})=\begin{cases}1,&\text{if }\text{complexity}(O_{j}^{(d)})>\theta\\ 0,&\text{otherwise}\end{cases}$ (3)
where $\text{complexity}(O_{j}^{(d)})$ denotes the complexity of direction $O_{j}^{(d)}$, and $\theta$ is the complexity threshold. In this work, we use a simple yet effective strategy in which each direction is treated as a section, and complexity is measured by the number of words.
For outline nodes requiring further decomposition, where $\phi(O_{j}^{(d)})=1$, the process is applied recursively as follows:
$\{L^{(d+1)}_{1},L^{(d+1)}_{2},\dots,L^{(d+1)}_{I}\}$ $=A_{S}(O_{j}^{(d)})$ (4)
$\{O_{1}^{(d+1)},O_{2}^{(d+1)},\ldots,O_{\mathcal{J}}^{(d+1)}\}$ $=A_{O}(\{L_{i}^{(d+1)}\})$
where $\{L_{i}^{(d+1)}\}$ denotes the set of literature retrieved for sub-direction $O_{j}^{(d)}$ at depth $d+1$, and $\mathcal{J}$ denotes the number of sub-directions.
This process constructs child nodes in the tree $\mathcal{T}$, where nodes $\{O_{1}^{(d+1)},O_{2}^{(d+1)},\ldots,O_{\mathcal{J}}^{(d+1)}\}$ are the children of node $O_{j}^{(d)}$. The exploration process terminates when the following condition is met:
$\forall v\in\mathcal{V}:\phi(v)=0\quad\text{or}\quad d=d_{\max}$ (5)
where $\mathcal{V}$ denotes all leaf nodes in the outline tree and $d_{\max}$ is the maximum allowed depth. The resulting tree $\mathcal{T}$ provides a hierarchical outline, with associated literature at each level, for downstream generation.
3.2 Exploitation Taskforce
The exploitation taskforce is designed to extract relevant snippets from the literature and draft the review manuscript. Without clear target content, it is difficult to locate the necessary evidence in the literature. Conversely, without sufficient supporting evidence, the manuscript may lack grounding. To mitigate compounding errors, we propose a tree-based strategy.
In detail, given the outline tree $\mathcal{T}$ produced as mentioned above, the manager $A_{M}$ initiates the exploitation taskforce. For each leaf node in the tree, $A_{M}$ identifies its parent outline node, denoted as $\{\mathcal{O}_{j}\}$, and the associated literature set $\{\mathcal{L}_{i}\}$. The exploitation taskforce then proceeds through iterative cycles. At the initial iteration $t=0$, the locating agent $A_{L}$ extracts relevant information based on the outline node $\{\mathcal{O}_{j}\}$ and the literature set $\{\mathcal{L}_{i}\}$, and operates as follows:
$\{F_{1}^{(0)},F_{2}^{(0)},\ldots,F_{N}^{(0)}\}=A_{L}(\{\mathcal{L}_{i}\},\{\mathcal{O}_{j}\})$ (6)
where $F_{n}^{(0)}$ denotes the $n$-th note snippet extracted from the literature set $\{\mathcal{L}_{i}\}$ relevant to the parent outline $\{\mathcal{O}_{j}\}$, and $N$ is the total number of snippets.
The drafting agent $A_{W}$ composes the initial draft as follows:
$M^{(0)}_{k}=A_{W}(\{F^{(0)}_{n}\},\{\mathcal{O}_{j}\})$ (7)
where $M^{(0)}_{k}$ represents the initial manuscript for the outline set $\{\mathcal{O}_{j}\}$, incorporating the content of all its child outline nodes.
For subsequent iterations $t>0$, the locating agent $A_{L}$ re-examines the literature to extract new snippets that support or improve the current draft, operating as follows:
$\{F_{1}^{(t)},F_{2}^{(t)},\ldots,F_{N}^{(t)}\}=A_{L}(\{\mathcal{L}_{i}\},M^{(t)}_{k})$ (8)
where $M^{(t)}_{k}$ denotes the draft content from the previous iteration, and $\{F^{(t)}_{n}\}$ is the set of newly extracted note snippets. The drafting agent $A_{W}$ refines the section draft by incorporating the new evidence and the previous draft:
$M^{(t+1)}_{k}=A_{W}(\{F^{(t)}_{n}\},\{\mathcal{O}_{j}\},M^{(t)}_{k})$ (9)
where $M^{(t+1)}_{k}$ is the updated content, and $M^{(t)}_{k}$ serves as the previous context for refinement.
The exploitation process for each parent node $\mathcal{O}_{i}$ terminates when the following condition is met:
$\psi(M^{(t+1)}_{i},M^{(t)}_{i})>\epsilon\quad\text{or}\quad t=t_{\max}$ (10)
where $\psi(\cdot,\cdot)$ denotes the similarity between consecutive iterations of the draft, measured by the ROUGE-1 score Lin (2004), $\epsilon$ is a similarity threshold that indicates no further significant improvements can be made, and $t_{\max}$ is the maximum number of allowed iterations.
The final content $M^{(t^{*})}_{i}$, where $t^{*}$ denotes the stopping iteration for parent node $i$, serves as the final draft. Finally, the manager $A_{M}$ aggregates the finalized drafts for all sections and delivers the completed literature review to the user.
3.3 Feedback Taskforce
The feedback taskforce is designed to execute tasks assigned by the manager, taking into account the risk of deviations at each step arising from unpredictable agent-environment interactions. To mitigate compounding errors, we propose an experience-based strategy.
In detail, for each agent $A_{W}$, the manager agent $A_{M}$ initiates the feedback taskforce. The agent first generates the primary results, then $A_{M}$ leverages historical experience to review these results, after which $A_{W}$ refines its outputs based on the feedback provided. At the initial step, $A_{W}$ produces the first output solely based on the context:
$Y^{(0)}=A_{W}(C)$ (11)
where $Y^{(0)}$ is the initial output, and $C$ is the context fetched by the manager agent.
[…]
Table 1: Comparison of automatic survey generation methods across different survey lengths (measured in tokens) on the benchmark dataset. Higher scores indicate better performance.
To give useful feedback, we propose an experience-aware mechanism. We record feedback and revision history in a database and select useful feedback for demonstration to provide feedback on the worker’s output at each iteration $x$.
$r^{*}=\arg\max_{r\in\mathcal{R}_{\min}}\text{usage}(r)$ (12)
where useful feedback is defined as the feedback that has led to the fewest further revisions, where $\mathcal{R}_{\min}=\{r\in\mathcal{R}:\text{revision}(r)=r_{\min}\}$. Here, $\text{revision}(r)$ denotes the number of revisions after using feedback $r$. In $\mathcal{R}_{\min}$, we select the most useful one, i.e., with the highest $\text{usage}(r)$. Here, $\text{usage}(r)$ denotes the number of times the feedback $r$ is selected for demonstration.
Referring to $r^{*}$, the manager agent then provides feedback on the worker’s output:
$R^{(x)}=A_{M}(Y^{(x)},r^{*})$ (13)
where $R^{(x)}$ denotes the feedback at iteration $x$.
For subsequent iterations ($x\geq 1$), the worker agent refines the output with the input and the feedback.
$Y^{(x)}=A_{W}(C,R^{(x-1)})$ (14)
The feedback taskforce continues until the output meets the manager’s requirements and no further feedback is provided, or until a maximum number of iterations $x_{\max}$ is reached. In the following, we detail how workers cooperate with the manager.
The searching agent formulates search queries based on user instructions, submits queries to search engines, and collects relevant literature. It filters and ranks the retrieved references by citation count, removes incomplete entries, and selects the top $Q_{k}$ unique references for further use. The outlining agent utilizes the gathered literature and user instructions to prompt an LLM for generating sub-outlines. The locating agent downloads and parses full-text papers, applies LLM-based previewing to select pertinent pages, and extracts concise text snippets for drafting. The drafting agent prompts the LLM to synthesize coherent text by incorporating selected references and prior drafts.
4 Experiments
4.1 Experimental Settings
We conduct experiments on two benchmark datasets, AutoSurvey Wang et al. (2024b) and SurveyEval Wang et al. (2025b), as well as a self-constructed benchmark dataset, TopSurvey. For hyperparameters, we set $d_{\max}$, $t_{\max}$, and $x_{\max}$ to 4. The hyperparameter $\theta$ is set to 500. We use GPT-4.1 for all agents. The evaluation comprises two categories of metrics. For citation quality, we adopt citation recall and citation precision as proposed by Wang et al. (2024b): recall measures whether cited passages fully support all statements, while precision measures the proportion of relevant citations that support their corresponding statements. For content quality, following Wang et al. (2024b), we use coverage, structure, and relevance, each rated by LLMs on a 5-point scale. Coverage assesses the extent to which the survey addresses all relevant aspects of the topic; structure evaluates logical organization and coherence; and relevance measures alignment with the specified research topic. We do not filter out fractional scores, such as 4.5.
4.2 Comparision Experiments
Figure 3: Comparison of automatic survey generation methods on the SurveyEval benchmark. Higher scores indicate better performance.
We first evaluate our framework on the AutoSurvey Wang et al. (2024b) benchmark, following the protocols established by Liang et al. (2025); Wang et al. (2024b). We use the same 20 topics from diverse subfields of LLM research to generate survey articles for comparison. We compare our approach with three baselines: Naive RAG-based LLMs using retrieval-augmented generation, AutoSurvey Wang et al. (2024b), and SurveyX Liang et al. (2025). We follow the official report of AutoSurvey [^1] https://github.com/AutoSurveys/AutoSurvey and conduct experiments with the official implementation of the offline version’s SurveyX [^2] https://github.com/IAAR-Shanghai/SurveyX.
As shown in Table 1, most algorithms exhibit reduced performance as survey length increases, particularly in recall, precision, structure, and relevance. This decline is most pronounced in the Naive RAG baseline, indicating that longer workflows lead to greater error accumulation and propagation. In contrast, coverage remains stable or improves with longer surveys, likely due to more comprehensive topic inclusion. State-of-the-art methods such as AutoSurvey and SurveyX continue to face challenges with citation precision, frequently generating statements unsupported by references. For content quality, structure consistently receives the lowest scores, reflecting disorganized article organization. Our proposed method remains robust to these issues and consistently achieves superior results across all metrics.
We further evaluate our framework on the SurveyEval benchmark Wang et al. (2025b), using the same protocols. SurveyEval is the first benchmark in computer science that pairs surveys with complete reference papers, comprising 384 arXiv cs.CL surveys citing over 26,000 references. Twenty topics were selected for testing based on reference completeness and reference list diversity. Experimental results are shown in Fig. 3, with Coverage, Structure, and Relevance scores normalized to a 100-point scale.
[…]
Table 2: Ablation study of the proposed modules on the benchmark dataset.
In Fig. 3, all methods achieve high coverage and relevance, indicating that LLMs like GPT-4.1 can generate comprehensive and relevant content. However, recall and precision remain low, reflecting poor reference retrieval and insufficient support for generated statements. As a result, state-of-the-art methods struggle with logical organization, leading to lower structure scores. Our method overcomes these issues, achieving the best performance across all metrics.
[…]
Table 3: Comparison of automatic survey generation methods at a survey length of 64k tokens on the new large-scale benchmark dataset. Higher scores indicate better performance.
[…]
Table 4: Sensitivity experiments across different iterations for the experience, exploration, and exploitation taskforces on the new large-scale benchmark.
4.3 Ablation Study
We further conduct an ablation study to evaluate the effectiveness of each component in our proposed framework. Experiments are performed on the same benchmark dataset as before, with all experimental settings unchanged and the survey length set to 8k tokens. We compare the full model with three variants. We use a single round of retrieval and organization for the variant without the exploration taskforce to generate the overall outline. We draft the manuscript using only one round of extraction and writing for the variant without the exploitation taskforce. For the variant without the feedback taskforce, each agent completes its assigned task, without revision.
Table 2 presents the results. Removing the exploration taskforce causes the largest drops in recall and structure, demonstrating its importance for citation recall and logical organization. Excluding the exploitation taskforce sharply reduces precision and relevance, confirming its role in improving citation precision and content relevance. Without the feedback taskforce, all metrics decrease, especially coverage, highlighting its key role in ensuring comprehensive coverage through collaborative revision.
4.4 A New Large-Scale Benchmark
To further validate the effectiveness of our proposed framework, we construct a new large-scale benchmark dataset. This dataset consists of 195 topics from various subfields of computer science, nearly 10 times larger than previous benchmarks Wang et al. (2024b); Liang et al. (2025). To ensure topic quality, we collect peer-reviewed survey topics from top computer science conferences, rather than from preprint sources such as arXiv. Survey papers accepted only as abstracts are excluded. To prevent data leakage from LLM pretraining data, we include only survey papers published in 2023, 2024, and 2025. We employ PhD students in computer science to verify whether a paper qualifies as a survey and to collect the final set of 195 survey papers: 9 from AAAI, 34 from ACL, 46 from EMNLP, 3 from ICLR, 2 from ICML, 77 from IJCAI, and 24 from NAACL. Of these, 68 were published in 2023, 105 in 2024, and 22 in 2025.
We conduct experiments using the same settings as above, generating 64k-token literature reviews for evaluation. As shown in Table 3, compared with results on the existing benchmark in Table 1, performance on the new large-scale benchmark drops significantly, particularly in recall and coverage. This may be due to the greater number and broader range of topics, which leads to some relevant literature not being retrieved, resulting in lower citation recall and coverage. Despite this, our proposed method achieves over 80% in citation scores and an average content quality score of 4.92.
To verify the consistency between LLM evaluation methods and human evaluation, we follow Wang et al. (2024b) to conduct a meta-evaluation. Human experts rank the generated surveys, and we compare these rankings with those generated by LLMs using Spearman’s rank correlation coefficient. The LLM evaluation achieves a correlation of 0.5117, suggesting LLMs align well with human preferences.
4.5 Sensitivity Analysis
To gain deeper insight into our framework, we conduct a sensitivity analysis to observe how $d_{\max}$, $t_{\max}$, and $x_{\max}$ affect performance. We evaluate the experience, exploration, and exploitation taskforces across different numbers of iterations, employing a controlled variable approach: when varying one hyperparameter, all others are held constant as previously described.
The results are presented in Table 4. We find that the second and third iterations yield the most significant improvements. This indicates that errors can accumulate at various steps in the workflow, and iterative refinement effectively reduces intermediate errors, thereby enhancing the quality of the final results. Both the Experience and exploitation taskforces contribute most to faithfulness; from the first to the fourth iteration, the number of references supporting claims in the final output increases by a large margin. Additionally, the Experience and exploration taskforces contribute notably to coverage and structural quality, suggesting that a single revision is often insufficient for comprehensive improvement. However, the effect of additional iterations gradually diminishes. By the fifth iteration, we observe a slight decline, due to unnecessary modifications overwriting correct results and introducing new errors.
Cluster Citation Quality Content Quality Recall $\uparrow$ Precision $\uparrow$ Coverage $\uparrow$ Structure $\uparrow$ Relevance $\uparrow$ 1 83.61 82.57 4.67 4.78 4.93 2 84.54 75.12 4.75 4.85 4.85 3 72.29 76.11 4.59 4.88 4.87 4 86.63 81.98 4.85 4.90 5.00 5 81.37 69.77 4.66 4.74 4.73
Table 5: Real-world case study across different topics on 400 generated results. Higher scores indicate better performance.
4.6 Disscussions
We evaluated our framework in a real-world setting by deploying an online automatic literature review generation system [^3] Anonymous for review; to be revealed upon acceptance., which has produced over 20,000 reviews. We randomly selected 400 reviews, embedded them with all-MiniLM-L6-v2, and applied K-means clustering to form five groups. Results show differences in citation and content quality among clusters. The best performance appears in computer science (cluster 4), while education and social sciences (cluster 3) perform worse, especially in citation quality. It may result from the search engine not including part of the relevant literature in the social sciences. We also measured system efficiency, and generating an 8,000-token review takes an average of 8.45 minutes.
5 Conclusion
In this paper, we identified and addressed the critical challenge of compounding errors in automated long-form literature review generation. The proposed Multi-Agent Taskforce Collaboration framework strategically orchestrates LLM-based agents through three specialized taskforces. Extensive experiments demonstrate that MATC achieves state-of-the-art performance across established benchmarks. The robustness and generalizability of our framework are further confirmed on our newly contributed TopSurvey benchmark, a large-scale collection of 195 peer-reviewed topics. The deployment of our online system, which has generated over 20,000 reviews, provides practical evidence of MATC’s effectiveness and efficiency in real-world scenarios.